OpenStudy (jennychan12):

Find the area of the region bounded by the graph of the equation. y = 3x^2 +1; x=0, x= 8, y =0

OpenStudy (amistre64):



OpenStudy (amistre64):

integrate y from 0 to 8

OpenStudy (jennychan12):

what's the y for then?

OpenStudy (amistre64):

y = f(x) \[\int_{a}^{b}f(x)~dx\]

OpenStudy (amistre64):

i spose if you wanted to subtract the bottom y function from the top y function .... but f(x)-0=f(x)

OpenStudy (amistre64):

in general, the area between 2 curves; say f(x) and g(x) is:\[\int_{a}^{b}f(x)-g(x)~dx\]

OpenStudy (jennychan12):

its ok, i know how to find the area.

OpenStudy (amistre64):

i knew you did :)

OpenStudy (jennychan12):

sorry, for another question. y = x^3+x, and x=2, and y = 0 how would i do that?

OpenStudy (amistre64):

those the "bounds"?

OpenStudy (jennychan12):

oh wait i think i know what it means. sorry.

OpenStudy (amistre64):